NEXT
Make New Entry,
Make Followup Entry
User name Sarah
Log entry time 00:10:30 on May 10, 2012
Entry number 383537
keyword=Swing Shift Summary
Swing Shift Summary
Leader: Sarah Phillips
Target: Mikhail Yurov
Third: Vincent Sulkosky
Summary
Most of the shift we had no beam, first because of beam studies, and then
because we were fixing the local dump viewer camera so that the beam
could be tuned to the local dump successfully. After the camera was
fixed and we went to beam permit, the accelerator people were able to aim
the beam to the center of the local dump, as confirmed with the camera,
within seconds. Then we went to the BPM calibration plan.
Timeline
- 16:00 - Shift starts with no beam, with the hall is restricted
access while the accelerator people are doing the beam studies. The
target people are finishing up work on the target, Jack is looking at the
septa, Ed is working on the dump viewer camera, and a halog entry tells
me that the sieves are in (halog
383480).
- 17:40 - The sieve is now out (halog
383500 ).
- 18:15 - Vince has ramped the other magnets, and is now ramping the
septa.
- 17:56 - We lost the dump viewer camera again. Dustin called up from
the hall and said that he has to get on the platform, but to do that, the
septum magnets have to be ramped down and locked out, so I called Jack,
who is on his way in to do this.
- 20:40 - MCC called. They are ready to send beam. But we aren't!
Still no dump viewer camera. Jack, Dustin, and Hovhannes are working on
it.
- 21:24 - The dump viewer camera is up and running (see halog
383503 and halog
383504). I have requested that we have the hall swept so we can go
into controlled access.
- 22:05 - Karl's new run plan puts the right septum current over
Jack's guidelines (it would be at 888A instead of the 875 limit in Jack's
halog (halog
383492). Couldn't get Karl for confirmation, so Hovhannes called
Jack for permission, which was granted. However, we ramped the right
septum to 867 A for now instead of leaving it at 888 A for no reason.
We'll ramp it up when it is time to take data.
- 22: 29 - Hall A is in power permit.
- 22:34 - Hall A is in beam permit!
- 22:40 - MCC is really concerned about the fact that the "Hall A
overview screen" (the one with our bpms and locks and such displayed on
it - HallA_O_status.eld) reads that our magnet is at 41.08 Tesla, and
they don't seem to like that we say the value must be wrong. They say
they need that value to read back, but later they said that they didn't.
??
- 22:54 - beam in the hall, MCC is tuning.
- 23:20 - Hall C wanted to flip the IHWP, so it is now out!
- 23:24 - We can see the spot on the viewer, it is centered the first
time! Tah dah! We can also see it with a slow raster of 0.9 V.
- 23:32 - Things look good, so we are going to the BPM calibration
plan now. That starts with a harp scan, in tune mode at 5 muA.
- 23:38 - MCC calls; they have to do something in the injector that
will take 5 minutes (no beam in that time).
- 24:00 - End of shift! Beam in hall!
Notes for the next shift:
- IMPORTANT - The right septum is presently lower than needed for the
next measurement so it wasn't sitting at high field while we were tuning
up. Don't forget to ramp it up to the nominal current before going to
production data taking!
- Keep an eye on the energy value to make sure it doesn't wander.
- Hall C is now in control of the IHWP, so keep track of when they
flip it. Ideally, you want an ihwp flip about halfway through each
point. If they don't flip it for some reason, make sure that you do.
- Keep an eye on the beam current so it doesn't drift down to 40 nA.
Keep and eye on the beam positions too.
- If MCC cannot deliver beam due to an FSD on the Moller target when
you know the target is out, ask them to reboot the IOC.
Runs Taken
None.